### 07 Jan 2024 v5

@@ III-A Motivation
$H(\hat{\bm{y}})\geq H(\hat{\bm{y}}|\bm{ctx}),$ (3)
where $H$ denotes Shannon entropy, $\bm{ctx}$ is the context of $\hat{\bm{y}}$. Exploiting correlations in $\hat{\bm{y}}$ results in bit savings.
In Fig. 2 and Fig. 3, channel-wise correlations and spatial correlations in latent representation of Kodim19 extracted by Cheng’20 [24] are illustrated.
@@
However, existing entropy models fail to capture correlations in local spatial, global spatial, and channel domains. Spatial context modules have limited interactions between channels, while channel-wise context modules lack interaction within the current slice. Moreover, extending these models to high-resolution image coding with acceptable even linear complexity is non-trivial. These challenges, along with the potential to enhance rate-distortion performance, motivate us to design a linear complexity attention-based multi-reference entropy model. Our proposed linear complexity attention-based multi-reference entropy model effectively captures correlations in local spatial, global spatial, and channel domains, while maintaining a modest complexity for high-resolution image coding. Further details on our model are presented in the subsequent sections.
III-B Overall Architecture
III-B1 ${}^{++}$
The overall architecture of proposed model is illustrated in Fig. 9. This model is named ${}^{++}$ to distinguish it from MLIC, and ${}^{+}$, which are introduced in our conference version [30]. The architecture of ${}^{++}$, as depicted in Figure 1, incorporates the analysis transform $g_{a}$, synthesis transform $g_{s}$, hyper analysis $h_{a}$, and hyper synthesis $h_{s}$, which are simplified versions of Cheng’20 [24]. To reduce complexity, attention modules are removed.
@@ III-B1 ${}^{++}$
TABLE II: Settings of ${}^{++}$ and ${}^{++}$.
Figure 5: Linear Multi-Reference Entropy Model ${}^{++}$. The figure illustrates the process of decoding a slice $\hat{\bm{y}}^{i}$.
The hyper-parameters and settings of ${}^{++}$ are presented in Table II. Same to Minnen et al. [23], we adopt mixed quantization, which involves adding uniform noise for entropy estimation and utilizing STE [16] to ensure differentiability in the quantization process. Gaussian mean-scale distribution is adopted for entropy estimation. For latent representation $\bm{y}$, the quantization and estimated rate is formulated as:
$\hat{\bm{y}}=\mathrm{STE}(\bm{y}-\bm{\mu})+\bm{\mu},\$ (4)
@@
where $\bm{\mu}$ is the estimated mean of latent representation $\bm{y}$, $\bm{u}\sim\mathcal{U}(-0.5,0.5)$, $\mathcal{R}_{\hat{\bm{y}}}$ is the estimated rate of $\hat{\bm{y}}$.
III-B2 ${}^{++}$
The proposed linear complexity attention-based multi-reference entropy model effectively captures channel-wise, local spatial, and global spatial correlations with linear complexity. The linear complexity entropy model is denoted as ${}^{++}$ to distinguish it from MEM, and ${}^{+}$, which are proposed in our conference version [30]. To capture multi-correlations, the proposed ${}^{++}$ consists of four components: channel-wise context module $g_{ch}$, local spatial context module $g_{lc}$, intra-slice global spatial context module $g_{gc,intra}$, and inter-slice global spatial context module $g_{gc,inter}$. In the channel-wise context module, the latent representation $\hat{\bm{y}}$ is divided into slices $\{\hat{\bm{y}}^{0},\hat{\bm{y}}^{1},\cdots\,\hat{\bm{y}}^{L}\}$ [23] along the channel dimension, $L$ is the number of slices. For the $i$-th slice $\hat{\bm{y}}^{i}$, the channel-wise context module captures the channel-wise context $\bm{\Phi}_{ch}^{i}$ from slices $\hat{\bm{y}}^{<i}$. To capture local spatial correlations, checkerboard pattern [40] is employed, where the latent representation $\hat{\bm{y}}^{i}$ is divided into anchor part $\hat{\bm{y}}_{ac}^{i}$ and non-anchor part $\hat{\bm{y}}_{na}^{i}$. $\hat{\bm{y}}_{ac}^{i}$ is local-context-free. Local spatial context ${\bm{\Phi}}_{lc}^{i}$ of $\hat{\bm{y}}_{na}^{i}$ is captured from $\hat{\bm{y}}_{ac}^{i}$. We propose Overlapped Window-based Checkerboard Attention $g_{lc,attn}$ for better non-linearity and adaptability to capture local spatial contexts. The global contexts $\bm{\Phi}_{gc}$ of $i$-th slice are extracted from two dimensions: intra-slice contexts $\bm{\Phi}_{gc,intra}^{i}$, and inter-slice contexts $\bm{\Phi}_{gc,inter}^{i}$. We propose Intra-Slice Global Context Module $g_{gc,intra}^{i}$ and Inter-Slice Global Context Module $g_{gc,inter}$ to capture such correlations. Since different slices share the similar global similarity [30, 14], the global similarity of $\hat{\bm{y}}^{i-1}$ is employed to predict the global correlations between $\hat{\bm{y}}^{i}_{ac}$ and $\hat{\bm{y}}^{i}_{na}$. The inter-slice global context $\Phi^{i}_{gc,inter}$ are extracted from slices $\hat{\bm{y}}^{<i}$ via the global similarity of slices $\hat{\bm{y}}^{<i}$. We introduce these modules in the following sections. The structure of ${}^{++}$ is illustrated in Table II. We use Equation 1 as our loss function and the estimated rate can be formulated as: $\mathcal{R}=\mathcal{R}_{\hat{\bm{z}}}+\sum^{L}_{i=0}\left(\mathcal{R}_{\hat{\bm{y}}^{i}_{ac}}+\mathcal{R}_{\hat{\bm{y}}^{i}_{na}}\right)$, where
@@ III-B2 ${}^{++}$
$\mathcal{R}_{\hat{\bm{z}}}$ is the rate of side information, $\mathcal{R}_{\hat{\bm{y}}^{i}_{ac}}$ denotes the rate of the anchor part of $i$-th slice, $\mathcal{R}_{\hat{\bm{y}}^{i}_{na}}$ denotes the rate of the non-anchor part of $i$-th slice, $\bm{\Phi}_{h}$ is the hyper-priors extracted by hyper analysis $h_{a}$ and hyper synthesis $h_{s}$.
III-C Channel-wise Context Module
To extract channel-wise contexts, the latent representation $\hat{\bm{y}}$ is first evenly divided into multiple slices $\{\hat{\bm{y}}^{0},\hat{\bm{y}}^{1},\cdots,\hat{\bm{y}}^{L}\}$ along the channel dimension. Slice $\hat{\bm{y}}^{i}$ is conditioned on slices $\hat{\bm{y}}^{<i}$. A channel context module $g_{ch}$ is employed to squeeze and extract context information from $\hat{\bm{y}}^{<i}$ when encoding and decoding $\hat{\bm{y}}^{i}$. $g_{ch}$ consists of three $3\times 3$ convolutional layers. The channel context becomes ${\bm{{\bm{\Phi}}}}_{ch}^{i}=g_{ch}(\hat{\bm{y}}^{<i})$. The channel-wise context module $g_{ch}$ is able to refer to symbols in the same and close position in the previous slices and helps select the most relative channels and extract information beneficial for accurate probability estimation. The channel number of each slice $S$ is a hyper-parameter. Following Minnen et al [23], we set $S$ to $32$ and $L$ to $10$ in our model. Following existing methods [23, 25], latent residual prediction (LRP) modules [23] are adopted to predict quantization error according to decoded slices and hyper-priors $\bm{\Phi}_{h}$. Since the channel number of latent representation is freezed during training and inference and the number of slices is quite small, the encoding speed and decoding speed is still fast enough in spite of serial process among slices.
@@ III-C Channel-wise Context Module
III-D Checkerboard Attention-based Local Context Module
One limitation of CNN-based local context modules is their fixed weights, which restricts their ability to capture content-adaptive contexts. We argue that context-adaptation is essential due to the vast diversity of images. In transformers [43, 51, 44], the attention weight is generated dynamically according to the input, which inspires us to design a transformer-based content-adaptive local context module. The local receptive field can be envisioned as a window, where local spatial contexts are captured by dividing the feature map into windows. Since each symbol is most relevant to the symbols around it, we propose to make the divided windows overlapped. To achieve this, we propose the novel checkerboard attention context module $g_{lc,attn}$. The process of $i$-th slice is taken as an example. Assuming the resolution of the latent representation $\hat{\bm{y}}^{i}$ is $H\times W$, the stride is set to $1$ to divide $\hat{\bm{y}}^{i}$ into $H\times W$ overlapped windows and the window size is $K\times K$. To extract local correlations, the attention map of each window is computed at first. Same as the convolutional checkerboard context module, interactions between $\bm{y}_{ac}^{i}$ and $\bm{y}_{na}^{i}$ and interactions in $\bm{y}_{na}^{i}$ are not allowed. An example of the attention mask is illustrated in Fig. 7. Importantly, this attention mechanism does not alter the resolution of each window. Subsequently, A $K\times K$ convolutional layer is utilized to fuse local context information and and match the size of the local context with that of $\bm{y}^{i}$ before feeding it to a feed-forward network (FFN) [43]. The overall process is similar to standard transformer [43]. The process is formulated as:
${{\hat{\bm{y}}}^{i}}_{attn}=\textrm{softmax}\left(\frac{{{\hat{\bm{y}}}^{i}}_{ac,q}\times({{\hat{\bm{y}}}^{i}}_{ac,k})^{\top}}{\sqrt{S}}+mask\right)\times{{\hat{\bm{y}}}^{i}}_{ac,v},$ (9)
${{\hat{\bm{y}}}^{i}}_{conv}=\textrm{conv}_{K\times K}({{\hat{\bm{y}}}^{i}}_{attn}),$ (10)
${\bm{\Phi}}^{i}_{lc}=\textrm{FFN}({{\hat{\bm{y}}}^{i}}_{conv})+{{\hat{\bm{y}}}^{i}}_{conv},$ (11)
where $\hat{\bm{y}}^{i}_{ac,q},\hat{\bm{y}}^{i}_{ac,k},\hat{\bm{y}}^{i}_{ac,v}=\textrm{Embed}(\hat{\bm{y}}^{i}_{ac})$, $\hat{\bm{y}}^{i}_{ac}$ is anchor part of $i$-th slice, $mask$ the attention mask, $S$ is the channel number of each slice, FFN is the feed-forward neural network [43].
@@ III-D Checkerboard Attention-based Local Context Module
III-E Linear Complexity Intra-Slice Global Context Module
@@
Figure 7: Mask of Shifted Window-based Checkerboard Attention $g_{lc,attn}$. Blue squares are non-anchor part $\hat{\bm{y}}_{na}$, white squares are anchor part $\hat{\bm{y}}_{ac}$. Green square and yellow square are two windows.
Figure 8: Cosine similarity in the spatial domain of different slices of latent representation of Kodim19 extracted by Cheng’20 [24] (optimized for MSE, $\lambda=0.0483$) to visualize slices share similar global correlations.
@@
In latent representation $\hat{\bm{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 [24], as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\bm{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\bm{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\bm{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@
First, the vanilla approach is introduced. The process of $i-1$-th slice and the $i$-th slice are taken as an example. When compressing or decompressing $\hat{\bm{y}}^{i}$, the correlations between anchor part $\hat{\bm{y}}^{i-1}_{ac}\in\mathbb{R}^{L\times S}$ and non-anchor part $\hat{\bm{y}}^{i-1}_{na}\in\mathbb{R}^{L\times S}$ of slice $\hat{\bm{y}}^{i-1}$ are first computed. Because the checkerboard local context module makes anchor visible when decoding non-anchor part, we multiply the anchor part of current slice $\hat{\bm{y}}^{i}_{a}$ with the attention map between $\hat{\bm{y}}^{i-1}_{ac}$ and $\hat{\bm{y}}^{i-1}_{na}$, which is employed as the approximation of global similarities between $\hat{\bm{y}}^{i}_{ac}$ and $\hat{\bm{y}}^{i}_{na}$. Due to the local correlations, adjacent symbols have similar global correlations. A $K\times K$ convolutional layer is employed to refine the attention map by aggregating global similarities of adjacent symbols. The process of this Intra-Slice Global Context $g_{gc,inter}$ is parallel and is formulated as:
@@
$\hat{\bm{y}}^{i}_{attn}=\underbrace{\textrm{softmax}\left(\frac{{{\hat{\bm{y}}}^{i-1}}_{na,q}\times\left({{\hat{\bm{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)}_{\textrm{non-negative}}\times\hat{\bm{y}}^{i}_{ac,v},$ (12)
$\hat{\bm{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\bm{y}}^{i}_{attn})+\hat{\bm{y}}^{i}_{a,v},$ (13)
${\bm{\Phi}}^{i}_{gc,intra}=\textrm{DepthRB}(\hat{\bm{y}}^{i}_{conv}),$ (14)
@@
where $\hat{\bm{y}}^{i-1}_{na,q},\hat{\bm{y}}^{i-1}_{a,k}=\textrm{Embedding}\left(\hat{\bm{y}}^{i-1}\right)$, $\hat{\bm{y}}^{i}_{a,v}=\textrm{Embedding}\left(\hat{\bm{y}}^{i}_{a}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects [52, 53] of images. DepthRB is the depth-wise residual bottleneck [36] and is employed to enhance the non-linearity.

### 20 Feb 2024 v9

@@ 3.1 Motivation
$H(\hat{\boldsymbol{y}})\geq H(\hat{\boldsymbol{y}}|\boldsymbol{ctx}),$ (3)
where $H$ denotes Shannon entropy, $\boldsymbol{ctx}$ is the context of $\hat{\boldsymbol{y}}$. Exploiting correlations in $\hat{\boldsymbol{y}}$ results in bit savings.
In Fig. 2 and Fig. 3, channel-wise correlations and spatial correlations in latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) are illustrated.
@@
However, existing entropy models fail to capture correlations in local spatial, global spatial, and channel domains. Spatial context modules have limited interactions between channels, while channel-wise context modules lack interaction within the current slice. Moreover, extending these models to high-resolution image coding with acceptable even linear complexity is non-trivial. These challenges, along with the potential to enhance rate-distortion performance, motivate us to design a linear complexity multi-reference entropy model. Our proposed linear complexity multi-reference entropy model effectively captures correlations in local spatial, global spatial, and channel domains, while maintaining a modest complexity for high-resolution image coding. Further details on our model are presented in the subsequent sections.
3.2 Overall Architecture
3.2.1 ${}^{++}$
The overall architecture of proposed model is illustrated in Fig. 4. This model is named ${}^{++}$ to distinguish it from MLIC, and ${}^{+}$, which are introduced in our conference version (Jiang et al., 2023b). The architecture of ${}^{++}$, as depicted in Fig. 4, incorporates the analysis transform $g_{a}$, synthesis transform $g_{s}$, hyper analysis $h_{a}$, and hyper synthesis $h_{s}$, which are simplified versions of Cheng’20 (Cheng et al., 2020). To reduce complexity, attention modules are removed.
@@ 3.2.1 ${}^{++}$
Table 2: Settings of ${}^{++}$ and ${}^{++}$.
Figure 5: Linear Multi-Reference Entropy Model ${}^{++}$. The figure illustrates the process of decoding a slice $\hat{\boldsymbol{y}}^{i}$.
The hyper-parameters and settings of ${}^{++}$ are presented in Table 2. Same to Minnen et al. (Minnen & Singh, 2020), we adopt mixed quantization, which involves adding uniform noise for entropy estimation and utilizing STE (Theis et al., 2017) to ensure differentiability in the quantization process. Gaussian mean-scale distribution is adopted for entropy estimation. For latent representation $\boldsymbol{y}$, the quantization and estimated rate is formulated as:
$\hat{\boldsymbol{y}}=\mathrm{STE}(\boldsymbol{y}-\boldsymbol{\mu})+\boldsymbol{\mu},\$ (4)
@@
where $\boldsymbol{\mu}$ is the estimated mean of latent representation $\boldsymbol{y}$, $\boldsymbol{u}\sim\mathcal{U}(-0.5,0.5)$, $\mathcal{R}_{\hat{\boldsymbol{y}}}$ is the estimated rate of $\hat{\boldsymbol{y}}$.
3.2.2 ${}^{++}$
The proposed linear complexity multi-reference entropy model effectively captures channel-wise, local spatial, and global spatial correlations with linear complexity. The linear complexity entropy model is denoted as ${}^{++}$ to distinguish it from MEM, and ${}^{+}$, which are proposed in our conference version (Jiang et al., 2023b). To capture multi-correlations, the proposed ${}^{++}$ consists of four components: channel-wise context module $g_{ch}$, local spatial context module $g_{lc}$, intra-slice global spatial context module $g_{gc,intra}$, and inter-slice global spatial context module $g_{gc,inter}$. In the channel-wise context module, the latent representation $\hat{\boldsymbol{y}}$ is divided into slices $\{\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\cdots\,\hat{\boldsymbol{y}}^{L}\}$ (Minnen & Singh, 2020) along the channel dimension, $L$ is the number of slices. For the $i$-th slice $\hat{\boldsymbol{y}}^{i}$, the channel-wise context module captures the channel-wise context $\boldsymbol{\Phi}_{ch}^{i}$ from slices $\hat{\boldsymbol{y}}^{<i}$. To capture local spatial correlations, checkerboard pattern (He et al., 2021) is employed, where the latent representation $\hat{\boldsymbol{y}}^{i}$ is divided into anchor part $\hat{\boldsymbol{y}}_{ac}^{i}$ and non-anchor part $\hat{\boldsymbol{y}}_{na}^{i}$. $\hat{\boldsymbol{y}}_{ac}^{i}$ is local-context-free. Local spatial context ${\boldsymbol{\Phi}}_{lc}^{i}$ of $\hat{\boldsymbol{y}}_{na}^{i}$ is captured from $\hat{\boldsymbol{y}}_{ac}^{i}$. We propose Overlapped Window-based Checkerboard Attention $g_{lc,attn}$ for better non-linearity and adaptability to capture local spatial contexts. The global contexts $\boldsymbol{\Phi}_{gc}$ of $i$-th slice are extracted from two dimensions: intra-slice contexts $\boldsymbol{\Phi}_{gc,intra}^{i}$, and inter-slice contexts $\boldsymbol{\Phi}_{gc,inter}^{i}$. We propose Intra-Slice Global Context Module $g_{gc,intra}^{i}$ and Inter-Slice Global Context Module $g_{gc,inter}$ to capture such correlations. Since different slices share the similar global similarity (Jiang et al., 2023b; Guo et al., 2022), the global similarity of $\hat{\boldsymbol{y}}^{i-1}$ is employed to predict the global correlations between $\hat{\boldsymbol{y}}^{i}_{ac}$ and $\hat{\boldsymbol{y}}^{i}_{na}$. The inter-slice global context $\Phi^{i}_{gc,inter}$ are extracted from slices $\hat{\boldsymbol{y}}^{<i}$ via the global similarity of slices $\hat{\boldsymbol{y}}^{<i}$. We introduce these modules in the following sections. The structure of ${}^{++}$ is illustrated in Table 2. We use Equation 1 as our loss function and the estimated rate can be formulated as: $\mathcal{R}=\mathcal{R}_{\hat{\boldsymbol{z}}}+\sum^{L}_{i=0}\left(\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{ac}}+\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{na}}\right)$, where
@@ 3.2.2 ${}^{++}$
$\mathcal{R}_{\hat{\boldsymbol{z}}}$ is the rate of side information, $\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{ac}}$ denotes the rate of the anchor part of $i$-th slice, $\mathcal{R}_{\hat{\boldsymbol{y}}^{i}_{na}}$ denotes the rate of the non-anchor part of $i$-th slice, $\boldsymbol{\Phi}_{h}$ is the hyper-priors extracted by hyper analysis $h_{a}$ and hyper synthesis $h_{s}$.
3.3 Channel-wise Context Module
To extract channel-wise contexts, the latent representation $\hat{\boldsymbol{y}}$ is first evenly divided into multiple slices $\{\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\cdots,\hat{\boldsymbol{y}}^{L}\}$ along the channel dimension. Slice $\hat{\boldsymbol{y}}^{i}$ is conditioned on slices $\hat{\boldsymbol{y}}^{<i}$. A channel context module $g_{ch}$ is employed to squeeze and extract context information from $\hat{\boldsymbol{y}}^{<i}$ when encoding and decoding $\hat{\boldsymbol{y}}^{i}$. $g_{ch}$ consists of three $3\times 3$ convolutional layers. The channel context becomes ${\boldsymbol{{\boldsymbol{\Phi}}}}_{ch}^{i}=g_{ch}(\hat{\boldsymbol{y}}^{<i})$. The channel-wise context module $g_{ch}$ is able to refer to symbols in the same and close position in the previous slices and helps select the most relative channels and extract information beneficial for accurate probability estimation. The channel number of each slice $S$ is a hyper-parameter. Following Minnen et al (Minnen & Singh, 2020), we set $S$ to $32$ and $L$ to $10$ in our model. Following existing methods (Minnen & Singh, 2020; Zou et al., 2022), latent residual prediction (LRP) modules (Minnen & Singh, 2020) are adopted to predict quantization error according to decoded slices and hyper-priors $\boldsymbol{\Phi}_{h}$. Since the channel number of latent representation is freezed during training and inference and the number of slices is quite small, the encoding speed and decoding speed is still fast enough in spite of serial process among slices.
@@ 3.3 Channel-wise Context Module
3.4 Checkerboard Attention-based Local Context Module
One limitation of CNN-based local context modules is their fixed weights, which restricts their ability to capture content-adaptive contexts. We argue that context-adaptation is essential due to the vast diversity of images. In transformers (Vaswani et al., 2017; Dosovitskiy et al., 2021; Liu et al., 2021), the attention weight is generated dynamically according to the input, which inspires us to design a transformer-based content-adaptive local context module. The local receptive field can be envisioned as a window, where local spatial contexts are captured by dividing the feature map into windows. Since each symbol is most relevant to the symbols around it, we propose to make the divided windows overlapped. To achieve this, we propose the novel checkerboard attention context module $g_{lc,attn}$. The process of $i$-th slice is taken as an example. Assuming the resolution of the latent representation $\hat{\boldsymbol{y}}^{i}$ is $H\times W$, the stride is set to $1$ to divide $\hat{\boldsymbol{y}}^{i}$ into $H\times W$ overlapped windows and the window size is $K\times K$. To extract local correlations, the attention map of each window is computed at first. Same as the convolutional checkerboard context module, interactions between $\boldsymbol{y}_{ac}^{i}$ and $\boldsymbol{y}_{na}^{i}$ and interactions in $\boldsymbol{y}_{na}^{i}$ are not allowed. An example of the attention mask is illustrated in Fig. 7. Importantly, this attention mechanism does not alter the resolution of each window. Subsequently, A $K\times K$ convolutional layer is utilized to fuse local context information and and match the size of the local context with that of $\boldsymbol{y}^{i}$ before feeding it to a feed-forward network (FFN) (Vaswani et al., 2017). The overall process is similar to standard transformer (Vaswani et al., 2017). The process is formulated as:
${{\hat{\boldsymbol{y}}}^{i}}_{attn}=\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i}}_{ac,q}\times({{\hat{\boldsymbol{y}}}^{i}}_{ac,k})^{\top}}{\sqrt{S}}+mask\right)\times{{\hat{\boldsymbol{y}}}^{i}}_{ac,v},$ (9)
${{\hat{\boldsymbol{y}}}^{i}}_{conv}=\textrm{conv}_{K\times K}({{\hat{\boldsymbol{y}}}^{i}}_{attn}),$ (10)
${\boldsymbol{\Phi}}^{i}_{lc}=\textrm{FFN}({{\hat{\boldsymbol{y}}}^{i}}_{conv})+{{\hat{\boldsymbol{y}}}^{i}}_{conv},$ (11)
where $\hat{\boldsymbol{y}}^{i}_{ac,q},\hat{\boldsymbol{y}}^{i}_{ac,k},\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embed}(\hat{\boldsymbol{y}}^{i}_{ac})$, $\hat{\boldsymbol{y}}^{i}_{ac}$ is anchor part of $i$-th slice, $mask$ the attention mask, $S$ is the channel number of each slice, FFN is the feed-forward neural network (Vaswani et al., 2017).
@@ 3.4 Checkerboard Attention-based Local Context Module
3.5 Linear Complexity Intra-Slice Global Context Module
@@
Figure 7: Mask of Shifted Window-based Checkerboard Attention $g_{lc,attn}$. Blue squares are non-anchor part $\hat{\boldsymbol{y}}_{na}$, white squares are anchor part $\hat{\boldsymbol{y}}_{ac}$. Green square and yellow square are two windows.
Figure 8: Cosine similarity in the spatial domain of different slices of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize slices share similar global correlations.
@@
In latent representation $\hat{\boldsymbol{y}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, $C$ is the channel number, each channel contains distinct information, but they can be considered as thumbnails, as depicted in Fig. 2. Notably, the channels exhibit similar global similarities. This is evident from the visualization of cosine similarities between two slices of Cheng’20 (Cheng et al., 2020), as visualized in Fig. 8, where despite differences in magnitude, the global correlations are similar. When decoding the current slice $\hat{\boldsymbol{y}}^{i}\in\mathbb{R}^{L\times S}$, decoded slice $\hat{\boldsymbol{y}}^{i-1}\in\mathbb{R}^{L\times S}$ assists in estimating the global correlations in slice $\hat{\boldsymbol{y}}^{i}$. However, a challenge arises in determining how to estimate these global correlations. While cosine similarity may be useful, it is fixed and may not accurately capture the features. In this regard, attention maps prove to be a suitable choice. The embedding layer is learnable, which make it flexible to adjust the method for global correlations estimation by modifying queries, keys, and values.
@@
First, the vanilla approach is introduced. The process of $i-1$-th slice and the $i$-th slice are taken as an example. When compressing or decompressing $\hat{\boldsymbol{y}}^{i}$, the correlations between anchor part $\hat{\boldsymbol{y}}^{i-1}_{ac}\in\mathbb{R}^{L\times S}$ and non-anchor part $\hat{\boldsymbol{y}}^{i-1}_{na}\in\mathbb{R}^{L\times S}$ of slice $\hat{\boldsymbol{y}}^{i-1}$ are first computed. Because the checkerboard local context module makes anchor visible when decoding non-anchor part, we multiply the anchor part of current slice $\hat{\boldsymbol{y}}^{i}_{a}$ with the attention map between $\hat{\boldsymbol{y}}^{i-1}_{ac}$ and $\hat{\boldsymbol{y}}^{i-1}_{na}$, which is employed as the approximation of global similarities between $\hat{\boldsymbol{y}}^{i}_{ac}$ and $\hat{\boldsymbol{y}}^{i}_{na}$. Due to the local correlations, adjacent symbols have similar global correlations. A $K\times K$ convolutional layer is employed to refine the attention map by aggregating global similarities of adjacent symbols. The process of this Intra-Slice Global Context $g_{gc,inter}$ is parallel and is formulated as:
@@
$\hat{\boldsymbol{y}}^{i}_{attn}=\underbrace{\textrm{softmax}\left(\frac{{{\hat{\boldsymbol{y}}}^{i-1}}_{na,q}\times\left({{\hat{\boldsymbol{y}}}^{i-1}}_{ac,k}\right)^{\top}}{\sqrt{S}}\right)}_{\textrm{non-negative}}\times\hat{\boldsymbol{y}}^{i}_{ac,v},$ (12)
$\hat{\boldsymbol{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\boldsymbol{y}}^{i}_{attn}),$ (13)
${\boldsymbol{\Phi}}^{i}_{gc,intra}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$ (14)
@@
where $\hat{\boldsymbol{y}}^{i-1}_{na,q},\hat{\boldsymbol{y}}^{i-1}_{ac,k}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i-1}\right)$, $\hat{\boldsymbol{y}}^{i}_{ac,v}=\textrm{Embedding}\left(\hat{\boldsymbol{y}}^{i}_{ac}\right)$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. This is because the self attention is permutation-invariant and lacks inductive bias. Using a depth-wise convolutional for position embedding does have serval benefits. First, a $3\times 3$ depth-wise convolution is quite light, which has negligible influences on overall complexity. Second, a convolution-based position embedding is flexible for any resolution, due to its translation equivariance. Third, the convolution is able to embed position information because of the zero-padding and the boundary effects (Kayhan & Gemert, 2020; Islam et al., 2019) of images. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2023a) and is employed to enhance the non-linearity.